# Topological bounds on the dynamical growth rate of chemical reaction networks

### 03 Mar 2026 v1

@@ Abstract
Growth and decay are system-level properties of chemical reaction networks (CRNs) relevant from prebiotic chemistry to cellular metabolism. Their properties are typically analyzed through the kinetics of particular models, which requires specification of the full set of kinetic laws and parameters. In this work, we derive stoichiometry-based constraints on the growth (or shrinkage) rate, in the balanced-growth regime of scalable CRNs. The resulting bounds are controlled by a topological quantity, the maximum amplification factor, defined via a von Neumann max–min problem over feasible fluxes as illustrated by numerical tests on random-network ensembles of CRNs. We argue for the relevance of our results in the context of origin of life studies but also for designing synthetic chemical reaction networks.
Autocatalysis and positive feedback are ubiquitous in biochemical organization and in candidate prebiotic chemistries: they enable self-maintenance, replication-like amplification, and, in open settings, persistent growth [11, 4, 1]. Stoichiometric frameworks have recently made the identification of autocatalytic subnetworks operational via purely structural criteria and algorithms that identify minimal autocatalytic motifs and their embeddings inside large networks [6, 12, 16, 2]. This progress raises a natural quantitative question: once a network is topologically capable of amplification, what controls how fast it can grow or decay?
From a dynamical perspective, growth rates depend on nonlinear kinetics and on rate constants that can vary over orders of magnitude and are often poorly constrained. This creates a gap: topology-based notions of autocatalysis certify feasibility, while kinetics determines realized performance. Bridging this gap requires a mixed kinetic–topological metric that quantifies amplification strength. Prior structural results in mass-action settings exist in restricted regimes (e.g., special coupling or dilute limits) [8, 32, 24]; here we identify a broad balanced-growth regime in which a single metric controls the achievable growth rate.
@@
with the convention that species with $(\mathbb{S}^{-}\mathbf{x})_{s}=0$ do not constrain the minimum (they correspond to $\infty$). Note that $\mathbf{x}>\mathbf{0}$ denotes that $\mathbf{x}\geq\mathbf{0}$ but $\mathbf{x}\neq\mathbf{0}$. Inspired by von Neumann’s technological expansion rate in input-output models of economics ([25], [27], Sec. 9.5 in [14]), the MAF was introduced for hypergraphs in [5]. From now on, we will denote the MAF with $\alpha$ for clarity of presentation.
@@
For autonomous networks, it can be shown that the MAF satisfies $0<\alpha<\infty$ (SM). A CRN is autocatalytic if, in addition to being autonomous, there exists a reaction flux such that all species are net produced (SM). From Equation 3, it follows immediately that $\alpha>1$ for autocatalytic networks. In SM, we show that the MAF of reducible networks is determined by their irreducible component with the highest MAF; thus, $\alpha$ acts as a “best-module” topological score. For a given network, efficient algorithms and their implementation for identifying irreducible autocatalytic subnetworks with the highest MAF can be found in [5].
A network is considered to be minimal with respect to a certain property if it does not contain any subnetwork that also verify the property. In particular, finding the MAF for minimal autonomous networks reduces to solving the generalized-eigenvalue problem of the form $\mathbb{S}^{+}\mathbf{x}=\alpha\phantom{\cdot}\mathbb{S}^{-}\mathbf{x}.$ Minimal autocatalytic networks, termed (autocatalytic) cores, in reversible networks are shown to be of five distinct types [6], displayed in Fig. 1. These cores provide useful reference points for interpreting $\alpha$ as an amplification strength (see Table 1), yielding characteristic values such as $\sqrt{2}$ (Type I and III), the golden ratio $(\sqrt{5}-1)/2$ (Type IV) and $2$ (Type V). Numerical approximations of the exact values are summarized in Table 1. Since for minimal networks, Eq. 3 reduces to the generalized eigenvalue problem, the ‘dual’ autoinhibitory core resulting from reaction reversal exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ has MAF $1/\alpha$. Note that, however, for an arbitrary network, $\alpha(\mathbb{S}^{-},\mathbb{S}^{+})\neq 1/\alpha(\mathbb{S}^{+},\mathbb{S}^{-})$.
Figure 1: Five types of minimal autocatalytic cores (from Ref. [6]). Reversing all reactions exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ and maps the MAF as $\alpha\mapsto 1/\alpha$, yielding the corresponding dual autoinhibitory motif.
@@
Table 1: Representative MAF values for autocatalytic cores and their autoinhibitory duals, consistent with [6].
Balanced growth in scalable reaction networks— We consider the regime of balanced growth, which is typically used to study chemical attractors [29] and metabolic networks [10]. For species population vector $\mathbf{n}$, we introduce the extensive quantity $\Omega=\mathbf{w}^{T}\mathbf{n}$, where the vector $\mathbf{w}$ quantifies the relative weights of the species. Then, the relative population vector $\mathbf{q}=\mathbf{n}/\Omega$ is such that $\mathbf{w}^{T}\mathbf{q}=1.$ A choice of CRN kinetics $\mathbf{J}(\mathbf{y})$ is scalable [21] if it satisfies $\mathbf{J}(\mathbf{y})/c=\mathbf{J}(\mathbf{y}/c),$ for any $c>0$. For scalable fluxes, the rescaled reaction flux $\mathbf{J}(\mathbf{n})/\Omega=\mathbf{J}(\mathbf{q})$ only depends on $\mathbf{q}$. Under this condition, the composition dynamics of the relative population on the simplex reads (see SM for details):
@@
$\mathbf{J}_{r}(\mathbf{n})$ $=k_{r}\frac{1}{\Omega^{\|(\mathbb{S}^{-})^{r}\|-1}}\prod_{s\in\mathcal{S}}\mathbf{n}^{(\mathbb{S}^{-})^{r}_{s}},$ (6)
where $k_{r}>0$ is the rate constant associated with reaction $r$, $\|(\mathbb{S}^{-})^{r}\|=\sum_{s}(\mathbb{S}^{-})^{r}_{s}$ is the total order of the reactants. In contrast to standard mass-action kinetics, where $\Omega$ denotes a constant volume, which has a simple expression for ideal gas-phase systems [15], the choice $\Omega=\mathbf{w}^{T}\mathbf{n}$ is natural for modeling ideal fluids with $\mathbf{w}$ encoding the relative weights of the species contributing to the total volume.
@@
Numerical tests on random autonomous networks— We tested the bounds on randomly generated autonomous, unambiguous CRNs with $2\leq|\mathcal{S}|\leq 12$ and $2\leq|\mathcal{R}|\leq 12$ (SM). We also vary the maximum reaction order for products, $M^{+}$, and reactants, $M^{-}$, where $M^{\pm}:=\max_{r\in\mathcal{R}}\sum_{s\in\mathcal{S}}(\mathbb{S}^{\pm})_{s}^{r}.$ For each instance, we computed $\alpha$ from (3) using computational implementations in [27, 5]. When necessary, we also used integration of the scalable composition dynamics to its stable fixed point $\mathbf{q}^{*}$ following [21], and extracted $\Lambda$ from (5). Using $\delta\equiv\alpha-1$ and $\tilde{\Lambda}\equiv\Lambda/\|\mathbb{S}^{-}\|_{k}$, Equation 8 can be recast as
@@
Discussion and outlook— The bounds of Eq. (8) turn a purely structural quantity, the MAF $\alpha$, into a quantitative control parameter for balanced growth. Operationally, one can compute $\alpha$ from stoichiometry [5, 27, 7], then estimate $\|\mathbb{S}^{-}\|_{k}$ from input stoichiometries and rough kinetic scales, to obtain an upper bound on balanced growth.
@@
The second route is structural modification (adding/removing reactions, merging motifs, rewiring autocatalytic components), which changes $\alpha$ itself and thereby shifts the admissible bound. In evolutionary terms, this means a punctuated equilibrium with topological network changes playing the role of mutations that alter the admissible growth bound, while subsequent kinetic adaptation and neutral evolution adjust $\Lambda$ within the available range [26, 30, 17].
@@
Beyond this scenario, our result has also practical implications for open chemical reactors because in a continuously stirred reactor with dilution rate $d$, the population scale following a possible transient satisfies $\dot{\Omega}=(\Lambda-d)\,\Omega$. Then, sustained growth requires $\Lambda>d$, therefore $\Lambda$ acts as a dilution threshold above which the system is washed out. Combined with Eq. (8), this yields a necessary structural condition for persistence against dilution: even with optimal kinetic tuning, survival at dilution $d$ is excluded unless $\alpha>1$ and $d$ lies below the corresponding growth bound [18].
@@
In compartmentalized systems with a steady-composition, the growth rate $\Lambda$ controls the rate of change of any extensive quantity besides that of the global population [31]. The compartment volume is one such quantity, and indeed, experiments have shown that the compartment volume can inform on the autocatalytic growth of small confined chemical species and that it represents a proxy for fitness when coupling many such compartments together [22].
@@
These interpretations suggest several natural extensions of the present work. First, it opens a route towards network engineering: using only knowledge of stoichiometry, one could design networks, that could potentially drive growth above a target threshold, or induce a robust decay for $\alpha\leq 1$ (robust decay). Pursuing this program requires composition rules for $\alpha$: which elementary network operations (adding a reaction, merging species, etc.) increase or decrease $\alpha$ [28, 9], and by how much? For related results in the economics literature, see [3]. Such a set of design rules would make $\alpha$ predictive under network modification and enable sharp statements about minimal structural changes needed to shift admissible growth bounds.
@@
A second direction is to close the gap between topological capability ($\alpha$) and dynamical realization (the steady flux). Identifying sufficient dynamical conditions that force the steady flux to approach MAF-optimal patterns would turn the universal bound into a quantitative estimate for restricted kinetic classes and clarify which motifs control performance. At the same time, an improved understanding of the connection between the stoichiometric description of CRNs and thermodynamics [8] could uncover functional organization principles in CRNs energetics. While $\alpha$ is purely stoichiometric, sustained growth in open chemistry requires chemical work supplied by chemostats [20, 13, 23]. A natural extension is therefore to connect growth bounds to thermodynamic budgets stated in terms of chemostat chemical potentials, affinities, or dissipation rates.
@@
Finally, scalable balanced growth is an asymptotic statement. In finite populations, demographic noise and rare-event switching can induce extinctions or transitions between composition attractors even when a stable $\mathbf{q}^{*}$ exists deterministically [29, 19]. Extending the framework to stochastic scalable kinetics would quantify the reliability of a topologically permitted growth bounds and provide risk-aware structural criteria, such as expected persistence times under fluctuations.
@@
Acknowledgments— P.G. acknowledges Barnabé Ledoux, Eric Smith, Nathaniel Virgo, and Atsushi Kamimura for discussions. P.G. was funded by JST CREST (JPMJCR2011) and JSPS Grant No. 25H01365. J.W. acknowledges Chunqiu Xia for computational resources.
@@ A. Structural setting and definitions
$(\mathbb{S}^{-})_{sr}\geq 0,\qquad(\mathbb{S}^{+})_{sr}\geq 0,$ (S1)
@@
A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autonomous if [3, 1]:
@@ A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autocatalytic if:
Note that requiring all reactions to be unambiguous-i.e. no species are shared between the reactants and products of any reaction-recovers the corresponding definitions of autonomy and autocatalysis in [2]. The results in this work does not require this assumption.
@@
A CRN is reducible [4, 6] if there exists a semi-autonomous subnetwork inside the whole network, that is able to produce every species in itself without consuming species from outside. Mathematically, a CRN $(\mathcal{S},\mathcal{R})$ is reducible if there exists $(\mathcal{S}^{\prime},\mathcal{R}^{\prime})$ such that

### 30 Mar 2026 v2

@@ Abstract
Growth and decay are system-level properties of chemical reaction networks (CRNs) relevant from prebiotic chemistry to cellular metabolism. Their properties are typically analyzed through the kinetics of particular models, which requires specification of the full set of kinetic laws and parameters. In this work, we derive stoichiometry-based constraints on the growth (or shrinkage) rate, in the balanced-growth regime of scalable CRNs. The resulting bounds are controlled by a topological quantity, the maximum amplification factor, defined via a von Neumann max–min problem over feasible fluxes as illustrated by numerical tests on random-network ensembles of CRNs. We argue for the relevance of our results in the context of origin of life studies, but also for designing synthetic chemical reaction networks.
Autocatalysis and positive feedback are ubiquitous in biochemical organization and in candidate prebiotic chemistries: they enable self-maintenance, replication-like amplification, and, in open settings, persistent growth [eigen1977principle, baum2023ecology, ameta_self-reproduction_2021]. Stoichiometric frameworks have recently made the identification of autocatalytic subnetworks operational via purely structural criteria and algorithms that identify minimal autocatalytic motifs and their embeddings inside large networks [blokhuis2020universal, gagrani2024polyhedral, golnik2025enumeration, andersen_defining_2020]. This progress raises a natural quantitative question: once a network is topologically capable of amplification, what controls how fast it can grow or decay?
From a dynamical perspective, growth rates depend on nonlinear kinetics and on rate constants that can vary over orders of magnitude and are often poorly constrained. This creates a gap: topology-based notions of autocatalysis certify feasibility, while kinetics determines realized performance. Bridging this gap requires a mixed kinetic–topological metric that quantifies amplification strength. Prior structural results in mass-action settings exist in restricted regimes (e.g., special coupling or dilute limits) [despons_structural_2024, unterberger2022stoechiometric, nandan_p_autocatalytic_2025]; here we identify a broad balanced-growth regime in which a single metric controls the achievable growth rate.
@@
with the convention that species with $(\mathbb{S}^{-}\mathbf{x})_{s}=0$ do not constrain the minimum (they correspond to $\infty$). Note that $\mathbf{x}>\mathbf{0}$ denotes that $\mathbf{x}\geq\mathbf{0}$ but $\mathbf{x}\neq\mathbf{0}$. Inspired by von Neumann’s technological expansion rate in input-output models of economics ([neumann1945model], [sargent2019neumann], Sec. 9.5 in [gale1989theory]), the MAF was introduced for hypergraphs in [blanco2025identifyingselfamplifyinghypergraphstructures]. From now on, we will denote the MAF with $\alpha$ for clarity of presentation.
@@
For autonomous networks, it can be shown that the MAF satisfies $0<\alpha<\infty$ (SM). A CRN is autocatalytic if, in addition to being autonomous, there exists a reaction flux such that all species are net produced (SM). From Equation 3, it follows immediately that $\alpha>1$ for autocatalytic networks. In SM, we show that the MAF of reducible networks is determined by their irreducible component with the highest MAF; thus, $\alpha$ acts as a “best-module” topological score. For a given network, efficient algorithms and their implementation for identifying irreducible autocatalytic subnetworks with the highest MAF can be found in [blanco2025identifyingselfamplifyinghypergraphstructures].
A network is considered to be minimal with respect to a certain property if it does not contain any subnetwork that also verify the property. In particular, finding the MAF for minimal autonomous networks reduces to solving the generalized-eigenvalue problem of the form $\mathbb{S}^{+}\mathbf{x}=\alpha\phantom{\cdot}\mathbb{S}^{-}\mathbf{x}.$ Minimal autocatalytic networks, termed (autocatalytic) cores, in reversible networks are shown to be of five distinct types [blokhuis2020universal], displayed in Fig. 1. These cores provide useful reference points for interpreting $\alpha$ as an amplification strength (see Table 1), yielding characteristic values such as $\sqrt{2}$ (Type I and III), the golden ratio $(\sqrt{5}-1)/2$ (Type IV) and $2$ (Type V). Numerical approximations of the exact values are summarized in Table 1. Since for minimal networks, Eq. 3 reduces to the generalized eigenvalue problem, the ‘dual’ autoinhibitory core resulting from reaction reversal exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ has MAF $1/\alpha$. Note that, however, for an arbitrary network, $\alpha(\mathbb{S}^{-},\mathbb{S}^{+})\neq 1/\alpha(\mathbb{S}^{+},\mathbb{S}^{-})$.
Figure 1: Five types of minimal autocatalytic cores (from Ref. [blokhuis2020universal]). Reversing all reactions exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ and maps the MAF as $\alpha\mapsto 1/\alpha$, yielding the corresponding dual autoinhibitory motif.
@@
Table 1: Representative MAF values for autocatalytic cores and their autoinhibitory duals, consistent with [blokhuis2020universal].
Balanced growth in scalable reaction networks— We consider the regime of balanced growth, which is typically used to study chemical attractors [singh2023multistable] and metabolic networks [dourado_analytical_2020]. For species population vector $\mathbf{n}$, we introduce the extensive quantity $\Omega=\mathbf{w}^{T}\mathbf{n}$, where the vector $\mathbf{w}$ quantifies the relative weights of the species. Then, the relative population vector $\mathbf{q}=\mathbf{n}/\Omega$ is such that $\mathbf{w}^{T}\mathbf{q}=1.$ A choice of CRN kinetics $\mathbf{J}(\mathbf{y})$ is a Class-1 system [pandey2020exponential], or scalable [lin2020origin], if it satisfies $\mathbf{J}(\mathbf{y})/c=\mathbf{J}(\mathbf{y}/c),$ for any $c>0$. For scalable fluxes, the rescaled reaction flux $\mathbf{J}(\mathbf{n})/\Omega=\mathbf{J}(\mathbf{q})$ only depends on $\mathbf{q}$. Under this condition, the composition dynamics of the relative population on the simplex reads (see SM for details):
@@
$\mathbf{J}_{r}(\mathbf{n})$ $=k_{r}\frac{1}{\Omega^{\|(\mathbb{S}^{-})^{r}\|-1}}\prod_{s\in\mathcal{S}}\mathbf{n}_{s}^{(\mathbb{S}^{-})^{r}_{s}},$ (6)
where $k_{r}>0$ is the rate constant associated with reaction $r$, $\|(\mathbb{S}^{-})^{r}\|=\sum_{s}(\mathbb{S}^{-})^{r}_{s}$ is the total order of the reactants. In contrast to standard mass-action kinetics, where $\Omega$ denotes a constant volume, which has a simple expression for ideal gas-phase systems [gillespie1992rigorous], the choice $\Omega=\mathbf{w}^{T}\mathbf{n}$ is natural for modeling ideal fluids with $\mathbf{w}$ encoding the relative weights of the species contributing to the total volume.
@@
Numerical tests on random autonomous networks— We tested the bounds on randomly generated autonomous, unambiguous CRNs with $2\leq|\mathcal{S}|\leq 12$ and $2\leq|\mathcal{R}|\leq 12$ (SM). We also vary the maximum reaction order for products, $M^{+}$, and reactants, $M^{-}$, where $M^{\pm}:=\max_{r\in\mathcal{R}}\sum_{s\in\mathcal{S}}(\mathbb{S}^{\pm})_{s}^{r}.$ For each instance, we computed $\alpha$ from (3) using computational implementations in [sargent2019neumann, blanco2025identifyingselfamplifyinghypergraphstructures]. When necessary, we also used integration of the scalable composition dynamics to its stable fixed point $\mathbf{q}^{*}$ following [lin2020origin], and extracted $\Lambda$ from (5). Using $\delta\equiv\alpha-1$ and $\tilde{\Lambda}\equiv\Lambda/\|\mathbb{S}^{-}\|_{k}$, Equation 8 can be recast as
@@
Discussion and outlook— The bounds of Eq. (8) turn a purely structural quantity, the MAF $\alpha$, into a quantitative control parameter for balanced growth. Operationally, one can compute $\alpha$ from stoichiometry [blanco2025identifyingselfamplifyinghypergraphstructures, sargent2019neumann, de_martino_von_2012], then estimate $\|\mathbb{S}^{-}\|_{k}$ from input stoichiometries and rough kinetic scales, to obtain an upper bound on balanced growth.
@@
The second route is structural modification (adding/removing reactions, merging motifs, rewiring autocatalytic components), which changes $\alpha$ itself and thereby shifts the admissible bound. In evolutionary terms, this means a punctuated equilibrium with topological network changes playing the role of mutations that alter the admissible growth bound, while subsequent kinetic adaptation and neutral evolution adjust $\Lambda$ within the available range [qian_chemical_2010, sole_fundamental_2024, jain_model_2001].
@@
Beyond this scenario, our result has also practical implications for open chemical reactors because in a continuously stirred reactor with dilution rate $d$, the population scale following a possible transient satisfies $\dot{\Omega}=(\Lambda-d)\,\Omega$. Then, sustained growth requires $\Lambda>d$, therefore $\Lambda$ acts as a dilution threshold above which the system is washed out. Combined with Eq. (8), this yields a necessary structural condition for persistence against dilution: even with optimal kinetic tuning, survival at dilution $d$ is excluded unless $\alpha>1$ and $d$ lies below the corresponding growth bound [kolchinsky2025thermodynamics].
@@
In compartmentalized systems with a steady-composition, the growth rate $\Lambda$ controls the rate of change of any extensive quantity besides that of the global population [sughiyama2022chemical]. The compartment volume is one such quantity, and indeed, experiments have shown that the compartment volume can inform on the autocatalytic growth of small confined chemical species and that it represents a proxy for fitness when coupling many such compartments together [lu_small-molecule_2024].
@@
These interpretations suggest several natural extensions of the present work. First, it opens a route towards network engineering: using only knowledge of stoichiometry, one could design networks, that could potentially drive growth above a target threshold, or induce a robust decay for $\alpha\leq 1$ (robust decay). Pursuing this program requires composition rules for $\alpha$: which elementary network operations (adding a reaction, merging species, etc.) increase or decrease $\alpha$ [sarkar_design_2019, despons_stability_2025], and by how much? For related results in the economics literature, see [bapat1995perron]. Such a set of design rules would make $\alpha$ predictive under network modification and enable sharp statements about minimal structural changes needed to shift admissible growth bounds.
@@
A second direction is to close the gap between topological capability ($\alpha$) and dynamical realization (the steady flux). Identifying sufficient dynamical conditions that force the steady flux to approach MAF-optimal patterns would turn the universal bound into a quantitative estimate for restricted kinetic classes and clarify which motifs control performance. At the same time, an improved understanding of the connection between the stoichiometric description of CRNs and thermodynamics [despons_structural_2024] could uncover functional organization principles in CRNs energetics. While $\alpha$ is purely stoichiometric, sustained growth in open chemistry requires chemical work supplied by chemostats [liang2024thermodynamic, gagrani2025thermodynamic, marehalli_srinivas_thermodynamics_2024]. A natural extension is therefore to connect growth bounds to thermodynamic budgets stated in terms of chemostat chemical potentials, affinities, or dissipation rates.
@@
Finally, scalable balanced growth is an asymptotic statement. In finite populations, demographic noise and rare-event switching can induce extinctions or transitions between composition attractors even when a stable $\mathbf{q}^{*}$ exists deterministically [singh2023multistable, lazarescu_large_2019]. Extending the framework to stochastic scalable kinetics would quantify the reliability of a topologically permitted growth bounds and provide risk-aware structural criteria, such as expected persistence times under fluctuations.
@@
Acknowledgments— P.G. acknowledges Barnabé Ledoux, Eric Smith, Nathaniel Virgo, and Atsushi Kamimura for discussions. P.G. and D.L. acknowledge P.G. was funded by JST CREST (JPMJCR2011) and JSPS Grant No. 25H01365. J.W. acknowledges Chunqiu Xia for computational resources. This research was partially conducted while visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP) Thematic Program on Biological Information Processing (TP24BI).
@@ A. Structural setting and definitions
$(\mathbb{S}^{-})_{s}^{r}\geq 0,\qquad(\mathbb{S}^{+})_{s}^{r}\geq 0,$ (S1)
@@
A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autonomous if [gagrani2024polyhedral, blanco2025identifyingselfamplifyinghypergraphstructures]:
@@ A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autocatalytic if:
Note that requiring all reactions to be unambiguous-i.e. no species are shared between the reactants and products of any reaction-recovers the corresponding definitions of autonomy and autocatalysis in [blokhuis2020universal]. The results in this work does not require this assumption.
@@
A CRN is reducible [gale1989theory, sargent2019neumann] if there exists a semi-autonomous subnetwork inside the whole network, that is able to produce every species in itself without consuming species from outside. Mathematically, a CRN $(\mathcal{S},\mathcal{R})$ is reducible if there exists $(\mathcal{S}^{\prime},\mathcal{R}^{\prime})$ such that